# Concurrent Composition for Differentially Private Continual Mechanisms

### v2

@@ 1.2. Our contributions
Composition theorems for differential privacy have two main uses. One is to bound the accumulated privacy loss when DP systems are used repeatedly over time. Another is to facilitate the design of complex DP algorithms from simpler ones. The concurrent composition theorems for interactive mechanisms have already found an application in the privacy analysis of matching algorithms in local edge differential privacy model [4]. (This work also has a result in the continual observation model, but the privacy is only proven for oblivious adversaries.)
Our concurrent composition theorem allows for a simpler and more modular privacy analysis of continual mechanisms built out of other continual mechanisms. In particular, we give a new privacy proof for the continual monotone histogram mechanism of Henzinger, Sricharan, and Steiner [12]. This mechanism involves concurrent executions of a simpler continual histogram mechanism and multiple instantations of the Sparse Vector Technique mechanism, and the original privacy analysis was rather intricate even for oblivious adversaries (because the mechanism itself is adaptive in its queries and dataset updates to the building-block mechanisms). With our concurrent composition theorems (both Theorem 1.2 and Theorem 1.5), we obtain a much simpler privacy proof. Moreover, for the case of approximate DP, the original proof was only for oblivious adversaries, whereas our proof also applies to adaptive adversaries.
@@
General Conditions for Privacy Analysis of Continual Mechanisms.
In the above application, the adversary does not directly interact with the sub-mechanisms, but rather it is the “main” HSS mechanism $\mathcal{M}$ that issues messages to its sub-mechanisms, based on the messages it receives. (See Figure 4 for an illustration.) These messages are not necessarily identical to the messages that $\mathcal{M}$ receives. Instead based on the input that $\mathcal{M}$ receives from an analyst or an adversary (passed through $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(\cdot)$), it generates an input sequence for its sub-mechanisms $\mathcal{M}_{i}$ that, on neighboring inputs to $\mathcal{M}$, must be identical or neighboring for their respective verification functions $f_{i}$. Note that $\mathcal{M}$ may fix the set of sub-mechanisms at initialization or select these mechanisms adaptively over time. For example, in the application above, the continual monotone histogram mechanism maintains $d$ fixed binary counters while adaptively instantiating an unbounded number of SVT and Laplace mechanisms as need.
The privacy analysis of such mechanisms $\mathcal{M}$ is challenging as the creation of sub-mechanisms and their inputs is adaptive and dependent on each other. Moreover, in contrast to concurrent composition of interactive mechanisms, $\mathcal{M}$ has direct access to the dataset updates, which cannot be revealed without differentially private processing (which needs to be done by the sub-mechanisms). We give a set of conditions for $\mathcal{M}$ and prove a theorem that shows that under these conditions the privacy of $\mathcal{M}$ can directly be computed by applying our concurrent composition theorems to the continual sub-mechanisms. The general idea is that $\mathcal{M}$ should only use its input sequence to generate inputs for the differentially private sub-mechanisms and any other action of $\mathcal{M}$, such as creating new sub-mechanism or generating a response, should be taken based only on the privatized outputs of the sub-mechanisms. This will ensure that $\mathcal{M}$ does not use any “privacy budget” in addition to the privacy budget used by the sub-mechanisms. For example, if $\mathcal{M}$ executes $k$ continual sub-mechanisms and, upon receiving neighboring inputs, issues neighboring message sequences for the sub-mechanisms, then the privacy of $\mathcal{M}$ equals the privacy of the concurrent composition of the sub-mechanism, as computed by Theorem 1.2. The same holds for the concurrent $k$-sparse parallel composition of sub-mechanisms, as computed in Theorems 1.3 and  1.5.
@@
  1. (1)
    Concurrent Composition of Continual Mechanisms: In the concurrent composition of fixed continual mechanism, $k$ continual mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ are fixed at the beginning, and an adversary adaptively issues pairs of messages for them such that for each $\mathcal{M}_{i}$, the sequences formed by the first messages of all message pairs and the sequence formed by the second messages of all message pairs are neighboring. A secret bit $b\in\{0,1\}$ determines whether each mechanism receives the first or second sequence from each pair. The adversary’s goal is to guess $b$ using the answers of all $k$ mechanisms. Theorem 4.10, which is the formal statement of Theorem 1.2 above, analyzes the privacy of this composition.
@@
    Concurrent $k$-Sparse Parallel Composition of Continual Mechanisms: This composition generalizes the concurrent composition of continual mechanisms with $k$ fixed parameters by permitting the creation of an arbitrary number of continual mechanisms. As before, $k$ privacy parameters $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ are fixed at the beginning. The adversary adaptively creates a potentially unlimited number of continual mechanisms and issues pairs of messages to each mechanism such that for all but at most $k$ of these mechanisms, each message pair consists of identical messages. Let $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ denote the exceptional mechanisms, which may receive differing message sequences. For each $1\leq i\leq k$, the mechanism $\mathcal{M}_{i}$ must be $(\epsilon_{i},\delta_{i})$-DP, and the sequences of the first and second messages for $\mathcal{M}_{i}$ must be neighboring.
@@
    As in previous compositions, a secret bit $b$ determines whether each mechanism receives the first or second message from each pair, and the adversary seeks to guess $b$ based on the outputs. Theorem 1.5 analyzes this composition in the special case where $\delta_{1}=\dots=\delta_{k}=0$ and $\epsilon_{1}=\dots=\epsilon_{k}=\epsilon_{0}$. The general version for non-identical $\epsilon_{i}$ parameters is given in Theorem 5.2.
  3. (3)
@@
    Concurrent $k$-Sparse Parallel Composition of Interactive Mechanisms: In this composition, the adversary creates an unbounded number of interactive mechanisms over time, instead of continual ones. As before, $k$ privacy parameter pairs $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ are fixed in advance. Each time the adversary creates an interactive mechanism, they also select a pair of neighboring datasets for it. Instead of message pairs, the adversary adaptively sends (single) queries to each mechanism. A secret bit $b$ determines whether the mechanism is run on the first or second dataset of its respective pair. The adversary then tries to infer $b$ from the responses of all mechanisms. Theorem 1.3 analyzes the privacy of this composition when all $\mathcal{M}_{i}$ are $(\epsilon_{0},\delta_{0})$-DP. The more general case for mechanisms with possibly different $(\epsilon_{i},\delta_{i})$ parameters is stated in Theorem 5.5.
@@
The paper is organized as follows. Section 2 introduces all necessary notation, including non-interactive mechanisms, interactive mechanisms, and interactive post-processing mechanism. Section 3 defines continual mechanisms, identifiers $\mathcal{I}$, verification functions $f$, verifiers $\mathcal{V}$ and differential privacy for continual mechanisms. Concurrent composition of continual mechanisms is defined in Section 4 and our first concurrent composition theorem is proven. Section 5 introduces concurrent $k$-sparse parallel composition of continual mechanisms and interactive mechanisms, proves the theorems for this type of composition and shows the counterexample. In Section 6, we state the conditions and the theorem for analyzing the privacy of complex continual mechanisms that use non-interactive, interactive and continual sub-mechanisms. Finally, in Section 7, we apply our theorem to analyze the privacy of the continual histogram mechanism of [12]. Appendix A presents our extension of Lyu’s result [18] to infinite answer spaces and an unbounded number of rounds of interaction.
@@ 2.1. Non-interactive Mechanisms (NIM)
$\mathrm{RR}_{\epsilon,\delta}(b)=\begin{cases}(b,\top)&\quad\text{w.p. $\delta$}\\ (b,\bot)&\quad\text{w.p. $(1-\delta)\frac{e^{\epsilon}}{1+e^{\epsilon}}$}\\ (1-b,\bot)&\quad\text{w.p. $(1-\delta)\frac{1}{1+e^{\epsilon}}$}\end{cases}$
@@
A mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $(\epsilon,\delta)$-DP w.r.t. a neighbor relation $\sim$ if and only if for any $x_{0}\sim x_{1}$, there exists a randomized ”post-processor” $\mathcal{P}:\{0,1\}\times\{\top,\bot\}\to\mathcal{Y}$ such that for both $b\in\{0,1\}$, $\mathcal{M}(x_{b})$ is identically distributed to $\mathcal{P}(\mathrm{RR}_{\epsilon,\delta})$.
@@ 2.3. Interactive Post-Processing (IPM)
Let $\mathcal{M}$ be an IM $\mathcal{M}$ with an initial state space $S_{\mathcal{M}}^{\text{init}}$ and $\mathcal{P}$ be an IPM with a singleton initial state space $S_{\mathcal{P}}^{\text{init}}=\{s\}$. Let $\sim$ be a neighbor relation on $S_{\mathcal{M}}^{\text{init}}$. Define a binary relation $\sim_{*}$ on $S_{\mathcal{P}\!\circ^{*}\!\mathcal{M}}^{\text{init}}$ such that two initial states $(s_{\mathcal{M}}^{1},s)$ and $(s_{\mathcal{M}}^{2},s)$ satisfy $(s_{\mathcal{M}}^{1},s)\sim_{*}(s_{\mathcal{M}}^{2},s)$ if and only if $s_{\mathcal{M}}^{1}\sim s_{\mathcal{M}}^{2}$. If $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. the neighbor relation $\sim$ against adaptive adversaries, then $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ is also $(\epsilon,\delta)$-DP w.r.t. the neighbor relation $\sim_{*}$ against adaptive adversaries.  [^2] In [10], the definitions of IMs and IPMs are more restrictive. Specifically, for any IM $\mathcal{M}$, $S_{\mathcal{M}}=Q_{\mathcal{M}}=A_{\mathcal{M}}=\{0,1\}^{*}$, and for any IPM $\mathcal{P}$, $S_{\mathcal{P}}=M_{\mathcal{P}}=\{0,1\}^{*}$; However, their results do not rely on these restrictions, and the interactive post-processing lemma holds for arbitrary sets $S_{\mathcal{M}}$, $Q_{\mathcal{M}}$, $A_{\mathcal{M}}$, $S_{\mathcal{P}}$, and $M_{\mathcal{P}}$.

### v3

@@ 1.2. Our contributions
Composition theorems for differential privacy have two main uses. One is to bound the accumulated privacy loss when DP systems are used repeatedly over time. Another is to facilitate the design of complex DP algorithms from simpler ones. The concurrent composition theorems for interactive mechanisms have already found an application in the privacy analysis of matching algorithms in local edge differential privacy model [4]. (This work also has a result in the continual observation model, but the privacy is only proven for oblivious adversaries, which allows it to be proven without using a composition theorem for continual mechanisms.) Our concurrent composition theorem allows for a simpler and more modular privacy analysis of continual mechanisms built out of other continual mechanisms.
There are various applications, where a continual mechanism executes multiple differentially private mechanisms, such as the Sparse Vector Technique (SVT) and continual counters, as sub-mechanisms and concurrently interacts with them. Let $\mathcal{M}$ denote such a composite mechanism. Upon receiving an input, $\mathcal{M}$ may create new sub-mechanisms, interact multiple times with the existing ones, and then return an output. $\mathcal{M}$ generates inputs for its sub-mechanisms based on its own inputs and the previous outputs of all sub-mechanisms.(See Figure 4 for an illustration.) Analyzing the privacy of such mechanism designs is challenging because the creation and inputs of sub-mechanisms are interdependent and adaptive. Even if the inputs of $\mathcal{M}$ are non-adaptive, the inputs to its sub-mechanisms may become adaptive: Let $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ denote the sub-mechanisms of $\mathcal{M}$. If the inputs of $\mathcal{M}_{1}$ depend on the outputs of $\mathcal{M}_{2}$ and the inputs of $\mathcal{M}_{2}$ depend on the outputs of $\mathcal{M}_{1}$, then $\mathcal{M}_{1}$ receives inputs correlated with its own past outputs. Later, we see an example of this situation, where a histogram mechanism creates $d$ private continual counters at the initialization and an adaptive number of SVTs and Laplace mechanisms over time. In that design, the inputs of the counters depend on the output of the (active instance of the) SVT and vice versa.
In such modular designs, we model $\mathcal{M}$ as the post-processing of a specific concurrent composition of its (continual) sub-mechanisms. Since $\mathcal{M}$ accesses the raw (secret) data itself, it could in principle incur additional privacy loss beyond that of the concurrent composition of the sub-mechanisms. To rule out such privacy loss, we introduce a set of conditions on $\mathcal{M}$ and prove that, under these conditions, the privacy guarantee of $\mathcal{M}$ is exactly that of the concurrent composition of its sub-mechanisms. Intuitively, $\mathcal{M}$ must restrict its use of raw inputs solely to generating inputs for differentially private sub-mechanisms, while all other actions, such as creating new sub-mechanisms or determining output values, must depend only on the privatized outputs of the sub-mechanisms. This will ensure that $\mathcal{M}$ does not use any “privacy budget” in addition to the “privacy budget” used by the sub-mechanisms.
Continual mechanisms are typically presented using pseudocode (rather than as formal state machines exchanging messages, which are used by our conditions on $\mathcal{M}$), where the pseudocode receives an input, updates its variables, and returns an output. We therefore reformulate our conditions in this setting. We call a variable a decision variable if it is either the output of a private sub-mechanism or computed solely based on other decision variables. Our conditions are now as follows:
  • Destination condition: Decisions about whether to produce an output, instantiate a new sub-mechanism (and of which type), or interact with an existing one (and which one) must be based only on decision variables.
  • Response condition: The value of $\mathcal{M}$’s output must be determined solely by decision variables.
  • Mapping condition: Fixing the values of decision variables, neighboring inputs to $\mathcal{M}$ must be mapped to input sequences for the sub-mechanisms with a structure for which concurrent composition privacy is private (This condition is referred to as $f\to f^{\prime}$ property in Section 6).
We show that if these conditions are satisfied, then the privacy guarantee of $\mathcal{M}$ is the same as that of the underlying concurrent composition. For example, if $\mathcal{M}$ executes $k$ continual sub-mechanisms and ensures that its neighboring inputs yield neighboring input sequences for each sub-mechanism, then the privacy guarantee of $\mathcal{M}$ matches that of the concurrent composition of its sub-mechanisms as given in Theorem 1.2. The same conclusion holds for the concurrent $k$-sparse parallel composition of sub-mechanisms, as established in Theorem 1.3 and Corollary 1.7.
@@
An Algorithmic Application.
We give a new privacy proof for the continual monotone histogram mechanism of Henzinger, Sricharan, and Steiner (HSS) [13]. This mechanism involves concurrent executions of a simpler continual histogram mechanism and multiple instantiations of the Sparse Vector Technique mechanism, and the original privacy analysis was rather intricate even for oblivious adversaries (because the mechanism itself is adaptive in its queries and dataset updates to the building-block mechanisms). Combining our concurrent composition theorems (both Theorem 1.2 and Corollary 1.7), we obtain a much simpler privacy proof. Moreover, for the case of approximate DP, the original proof was only for oblivious adversaries, whereas our proof also applies to adaptive adversaries.
@@
(1) Concurrent Composition of Continual Mechanisms: In the concurrent composition of fixed continual mechanism, $k$ continual mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ are fixed at the beginning, and an adversary adaptively issues pairs of messages for them such that for each $\mathcal{M}_{i}$, the sequences formed by the first messages of all message pairs and the sequence formed by the second messages of all message pairs are neighboring. A secret bit $b\in\{0,1\}$ determines whether each mechanism receives the first or second sequence from each pair. The adversary’s goal is to guess $b$ using the answers of all $k$ mechanisms. Theorem 4.9, which is the formal statement of Theorem 1.2 above, analyzes the privacy of this composition.
@@
(2) Concurrent $k$-Sparse Parallel Composition of Continual Mechanisms: This composition generalizes the concurrent composition of continual mechanisms with $k$ fixed parameters by permitting the creation of an arbitrary number of continual mechanisms. As before, $k$ privacy parameters $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ are fixed at the beginning. The adversary adaptively creates a potentially unlimited number of continual mechanisms and issues pairs of messages to each mechanism such that for all but at most $k$ of these mechanisms, each message pair consists of identical messages. Let $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ denote the exceptional mechanisms, which may receive differing message sequences. For each $1\leq i\leq k$, the mechanism $\mathcal{M}_{i}$ must be $(\epsilon_{i},\delta_{i})$-DP, and the sequences of the first and second messages for $\mathcal{M}_{i}$ must be neighboring. We note the choice of mechanisms $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ is made adaptively over time. Specifically, the adversary decides whether to issue a pair of differing messages for a mechanism based on the previous answers of all mechanisms, including the outputs of that mechanism.
@@
As in previous compositions, a secret bit $b$ determines whether each mechanism receives the first or second message from each pair, and the adversary seeks to guess $b$ based on the outputs. Let $\mathcal{M}_{j}^{\prime}$ denote the $j$-th created mechanism, and let $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ be its privacy parameter. Theorem 5.4, informally stated in Theorem 1.5, analyzes the privacy of the concurrent $k$-sparse parallel composition of continual mechanisms when $\sum\delta_{j}^{\prime}$ is upper bounded by a predetermined value. Corollary 5.5, informally stated in Corollary 1.7, sets this upper bound to 0 and provides privacy guarantees for the concurrent $k$-sparse parallel composition of purely differentially private continual mechanisms.
@@
(3) Concurrent $k$-Sparse Parallel Composition of Interactive Mechanisms: In this composition, the adversary creates an unbounded number of interactive mechanisms over time, instead of continual ones. As before, $k$ privacy parameter pairs $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ are fixed in advance. Each time the adversary creates an interactive mechanism, they also select a pair of neighboring datasets for it. Instead of message pairs, the adversary adaptively sends (single) queries to each mechanism. A secret bit $b$ determines whether the mechanism is run on the first or second dataset of its respective pair. The adversary then tries to infer $b$ from the responses of all mechanisms. Theorem 1.3 analyzes the privacy of this composition when all $\mathcal{M}_{i}$ are $(\epsilon_{0},\delta_{0})$-DP. The more general case for mechanisms with possibly different $(\epsilon_{i},\delta_{i})$ parameters is stated in Theorem 5.15.
@@
The paper is organized as follows. Section 2 introduces all necessary notation, including non-interactive mechanisms, interactive mechanisms, and interactive post-processing mechanism. Section 3 defines continual mechanisms, identifiers $\mathcal{I}$, verification functions $f$, verifiers $\mathcal{V}$ and differential privacy for continual mechanisms. Concurrent composition of continual mechanisms is defined in Section 4 and our first concurrent composition theorem is proven. Section 5 introduces concurrent $k$-sparse parallel composition of continual mechanisms and interactive mechanisms, proves the theorems for this type of composition and shows the counterexample. In Section 6, we state the conditions and the theorem for analyzing the privacy of complex continual mechanisms that use non-interactive, interactive and continual sub-mechanisms. In Section 7, we apply our theorem to analyze the privacy of the continual histogram mechanism of [13]. Finally, in Section 8 and Section 9, we prove two technical intermediate lemmas, required for the previous results.
@@ 2.1. Non-interactive Mechanisms (NIM)
$\mathrm{RR}_{\epsilon,\delta}(b)=\begin{cases}(\top,b)&\quad\text{w.p. $\delta$}\\ (\bot,b)&\quad\text{w.p. $(1-\delta)\frac{e^{\epsilon}}{1+e^{\epsilon}}$}\\ (\bot,1-b)&\quad\text{w.p. $(1-\delta)\frac{1}{1+e^{\epsilon}}$}\end{cases}$
@@
A mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $(\epsilon,\delta)$-DP w.r.t. a neighbor relation $\sim$ if and only if for any $x_{0}\sim x_{1}$, there exists a randomized ”post-processor” $\mathcal{P}:\mathcal{Y}\to\{0,1\}\times\{\top,\bot\}$ such that for both $b\in\{0,1\}$, $\mathcal{M}(x_{b})$ is identically distributed to $\mathcal{P}(\mathrm{RR}_{\epsilon,\delta}(b))$.
@@ 2.2. Interactive Mechanisms (IM)
An example of an IM is the interactive randomized response mechanism, denoted $\mathrm{IRR}_{\epsilon,\delta}$, which will repeatedly be used in our proofs. Ghazi et al. [10] define this mechanism to initially flip a coin and determine whether to answer all subsequent queries truthfully or in a privacy-preserving manner. Our definition of interactive randomized response mechanism is different but shares the same intuition.
Recall that the standard randomized response mechanism $\mathrm{RR}_{\epsilon,\delta}$ is an NIM that outputs a pair from the space $\{\top,\bot\}\times\{0,1\}$. In contrast, we define the interactive variant $\mathrm{IRR}_{\epsilon,\delta}$ to output the pair in two separate steps: it first produces an element from $\{\top,\bot\}$, and later returns a bit in $\{0,1\}$, such that the joint distribution of the two outputs matches that of $\mathrm{RR}_{\epsilon,\delta}$. Specifically, $\mathrm{IRR}_{\epsilon,\delta}$ is an IM with a single query message $q^{*}$. Given an initial state $b\in\{0,1\}$, $\mathrm{IRR}_{\epsilon,\delta}(b)$ responds to the first query $q^{*}$ with an element in $\{\top,\bot\}$ and returns a bit in $\{0,1\}$ when receiving $q^{*}$ for the second time. $\mathrm{IRR}_{\epsilon,\delta}(b)$ halts the communication upon receiving $q^{*}$ for the third time. The formal definition is given below.
Definition 2.9 ($\mathrm{IRR}_{\epsilon,\delta}$).
For any $\epsilon\geq 0$ and $0\leq\delta\leq 1$, the interactive randomized response mechanism $\mathrm{IRR}_{\epsilon,\delta}$ is an IM with state space $S=\{\text{null}\}\cup\{0,1\}\cup\{\top,\bot\}\times\{0,1\}$, initial state space $S^{\text{init}}=\{0,1\}$, and query space $Q=\{q^{*}\}$. $\mathrm{IRR}_{\epsilon,\delta}$ is represented by a randomized function that maps a state $s\in S$ and query $q^{*}\in Q$ to a new state $s^{\prime}\in S$ and a message $m$ as follows:
  • If $s$ is a bit $b\in\{0,1\}$ (i.e., $q^{*}$ is received for the first time), the output $m$ is set to $\bot$ with probability $\delta$ and to $\top$ with probability $1-\delta$. The new state is then set to $s^{\prime}=(m,b)$.
  • If $s$ is a pair $(\tau,b)$ in $\{\top,\bot\}\times\{0,1\}$ (i.e., $q^{*}$ is received for the first time), $s^{\prime}$ is set to null and $m$ is determined as follows:
    • If $\tau=\bot$ (i.e., the first response was $\bot$), then $m=b$ with probability $1$.
    • If $\tau=\top$ (i.e., the first response was $\top$), $m=b$ with probability $\frac{e^{\epsilon}}{1+e^{\epsilon}}$ and m=1bm=1-b with probability $\frac{1}{1+e^{\epsilon}}$.
@@ 2.3. Interactive Post-Processing (IPM)
Let $\mathcal{M}$ be an IM $\mathcal{M}$ with an initial state space $S_{\mathcal{M}}^{\text{init}}$ and $\mathcal{P}$ be an IPM with a singleton initial state space $S_{\mathcal{P}}^{\text{init}}=\{s\}$. For every pair of initial states $s_{0},s_{1}\in S_{\mathcal{M}}^{\text{init}}$, the following holds: If for every adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{M}(s_{0}))$ and $\mathrm{View}(\mathcal{A},\mathcal{M}(s_{1}))$ are $(\epsilon,\delta)$-indistinguishable, then for every adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{P}\!\circ^{*}\!\mathcal{M}(s_{0}))$ and $\mathrm{View}(\mathcal{A},\mathcal{P}\!\circ^{*}\!\mathcal{M}(s_{1}))$ are $(\epsilon,\delta)$-indistinguishable.  [^5] In [11], the definitions of IMs and IPMs are more restrictive. Specifically, for any IM $\mathcal{M}$, $S_{\mathcal{M}}=Q_{\mathcal{M}}=A_{\mathcal{M}}=\{0,1\}^{*}$, and for any IPM $\mathcal{P}$, $S_{\mathcal{P}}=M_{\mathcal{P}}=\{0,1\}^{*}$; However, their results do not rely on these restrictions, and the interactive post-processing lemma holds for arbitrary sets $S_{\mathcal{M}}$, $Q_{\mathcal{M}}$, $A_{\mathcal{M}}$, $S_{\mathcal{P}}$, and $M_{\mathcal{P}}$.